# Uncertainty-Aware Critic Augmentation for Hierarchical Multi-Agent EV Charging Control

### v1

@@ A Reproducibility
For the design of the penalty mechanism, the penalty charge is typically two or three times the basic capacity rate [19]. We design the penalty system based on the amount of electricity exceeding the contracted capacity (the upper bound of instant electricity consumption)[7, 19], with detailed settings shown in Table 4. Let the maximum instant electricity load during the testing period be $L_{max}$. If $L_{max}<C_{contract}$, then the penalty is set to zero. Given a threshold coefficient $0<\delta<1$, the penalty is calculated by dividing the excess electricity into two tiers:
@@
(A.1) $\hskip 0.0pt\text{Penalty Cost}=2\cdot R_{base}\cdot\min(L_{max},\delta\cdot C_{contract})$
@@
The charging capacity ($\mathcal{P}^{\max}$) is set to 150 (kW), the EV battery capacity $C^{i}$ (Def.2) is 60 (kwh), and the charging efficiency is 95%. For training our HUCA model, we use 2,500 episodes, with a replay buffer capacity of 30,000 for each agent, and a batch size of 1,024. The penalty coefficient $\varphi$ (Eq. 3.2) weight factors $\kappa$ (Eq. 3.2), $\omega$ (Eq. 3.10), and the importance of the uncertainity term $\rho$ (Eq. 3.6) are set to 0.1, 0.1, 0.5, and 10, respectively.

### v3

@@ A Reproducibility
For the design of the penalty mechanism, the penalty charge is typically two or three times the basic capacity rate [18]. We design the penalty system based on the amount of electricity exceeding the contracted capacity (the upper bound of instant electricity consumption)[7, 18], with detailed settings shown in Table 4. Let the maximum instant electricity load during the testing period be $L_{max}$. If $L_{max}<C_{contract}$, then the penalty is set to zero. Given a threshold coefficient $10<\delta<$ 1, the penalty is calculated by dividing the excess electricity into two tiers:
@@
(A.1) Penalty $\text{Penalty Cost}=2\cdot R_{base}\cdot\min(L_{max},\delta\cdot C_{contract})$
@@
The charging capacity ($\mathcal{P}^{\max}$) is set to 150 (kW), the EV battery capacity $C^{i}$ (Def. 2) is 60 (kwh), and the charging efficiency is 95%. For training our HUCA model, we use 1,500 episodes, with a replay buffer capacity of 30,000 for each low-level agent, and a batch size of 1,024. The penalty coefficient $\varphi$ (Eq. 3.2) weight factors $\kappa$ (Eq. 3.2), $\omega$ (Eq. 3.10), and the importance of the uncertainty term $\rho$ (Eq. 3.6) are set to 0.1, 0.1, 0.5, and 10, respectively.